OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ?

OpenStudy (across):

12x+9

OpenStudy (anonymous):

thankes!

OpenStudy (anonymous):

f(g(x)) = f(3x+2) = 4(3x+2) +1 = 12x + 8 + 1 = 12x + 9